mpiJava status

JDK

1.1.6  1.2b3
green  native
1.0 p4 patched hang (allgather and bsend) same as 1.1.6 somtimes hang(range.java)
mpich  p4  error(MPI.init)  same as 1.1.6 O.K but test suite exit in abort.java
1.1.1 p4 patched —  —  —
globus  error(MPI.init)  —  —
sun mpi  error(MPI.init)  ?  error(MPI.init)
lam mpi  test suite hang  ?  ?